# Dynamic Gaussians Mesh: Consistent Mesh Reconstruction from Monocular Videos

### v2

@@ Abstract
Modern 3D engines and graphics pipelines require mesh as a memory-efficient representation, which allows efficient rendering, geometry processing, texture editing, and many other downstream operations. However, it is still highly difficult to obtain high-quality mesh in terms of structure and detail from monocular visual observations. The problem becomes even more challenging for dynamic scenes and objects. To this end, we introduce Dynamic Gaussians Mesh (DG-Mesh), a framework to reconstruct a high-fidelity and time-consistent mesh given a single monocular video. Our work leverages the recent advancement in 3D Gaussian Splatting to construct the mesh sequence with temporal consistency from a video. Building on top of this representation, DG-Mesh recovers high-quality meshes from the Gaussian points and can track the mesh vertices over time, which enables applications such as texture editing on dynamic objects. We introduce the Gaussian-Mesh Anchoring, which encourages evenly distributed Gaussians, resulting better mesh reconstruction through mesh-guided densification and pruning on the deformed Gaussians. By applying cycle-consistent deformation between the canonical and the deformed space, we can project the anchored Gaussian back to the canonical space and optimize Gaussians across all time frames. During the evaluation on different datasets, DG-Mesh provides significantly better mesh reconstruction and rendering than baselines.
Keywords:
Dynamic Scene Reconstruction Surface Reconstruction Neural Rendering
footnotetext: Project page: https://www.liuisabella.com/DG-Mesh/footnotetext: Equal advisory
Figure 1: We propose DG-Mesh, a framework that reconstructs high-fidelity time-consistent mesh from monocular videos. Given monocular inputs of dynamic scenes and the camera parameters, our method reconstructs the high-quality surface and its appearance, as well as the mesh vertice motion across time frames. Our method can reconstruct mesh with flexible topology change. We can reconstruct satisfying surfaces even in challenging thin structures like bird wings.
@@ 1 Introduction
The advent of neural rendering techniques has sparked a surge in research focused on extracting the geometry and motion of dynamic scenes from monocular videos using neural representations. Most studies have concentrated on learning deformable neural radiance fields using volumetric representations [21]. However, these volumetric models often fall short in terms of memory efficiency and explicit geometric details. The recent introduction of 3D Gaussian Splatting [13] shifts towards point cloud representations that not only are more memory-efficient but also can provide explicit geometries and superior rendering quality. Further advancements demonstrate how 3D Gaussian Splatting’s explicit geometry can effectively represent dynamic scenes by tracking each point’s movement across frames [20, 39, 49].
This paper takes a significant step in the realm of explicit geometry and motion estimation for dynamic scenes captured as a monocular video. We introduce a method to extract high-fidelity meshes and their motions by tracking vertices over time from a monocular video, as illustrated in Fig. 3. Meshes, in contrast to volume, offer a more memory-efficient format. The mapping of correspondences on mesh vertices greatly simplifies texture editing and propagation on target objects. Crucially, the dynamic mesh extracted via our method can be seamlessly integrated into a physical simulator, enabling rapid, physics-based rendering adaptable to diverse materials and lighting conditions.
@@
We achieve this by decomposing the mesh-to-mesh correspondence into two separate correspondences that are easier to acquire: the mesh-to-points correspondence that maps the mesh faces to the Gaussian points in each time frame; and the points-to-canonical-points correspondence that helps tracking the movement of mesh across time. Our key observation is that the direct training of 3D Gaussians mentioned above will lead to uneven spreading of Gaussians in 3D space (as shown in Fig. 2 left). However, this is detrimental for mesh reconstruction performance due to the violation of the Possion solver assumption that points should be evenly distributed; worse still, non-uniform Gaussians make it trickier to track across a long time period. Thus, we propose the Gaussian-Mesh Anchoring procedure during training, which yields uniformly distributed Gaussians at each frame. Concretely, first, we encourage deformed 3D Gaussians in each time step to align with the corresponding mesh surface that has uniform face distribution due to the underlying iso-surface algorithm; and second, we allow adding Gaussians in the world frame at each time step if certain mesh faces are not covered by existing deformed Gaussians, as well as deduplicting Gaussians for mesh faces that are covered multiple times. We call the new set of Gaussians with 1-1 correspondences to mesh faces as anchored Gaussians. As depicted in Fig. 2, anchored Gaussians spread more uniformly in 3D.
Figure 2: 3D Gaussian points before and after applying our Gaussian-Mesh Anchoring.
@@
Finally, we need to update the canonical space to accommodate the added and removed Gaussians in the world frame so that canonical and anchored Gaussians are always consistent. This can be done by introducing a backward deformation network that computes the deformation from the anchored Gaussians to the canonical Gaussians, by which we can inject new Gaussians in the canonical space or remove unnecessary Gaussians. We define a Cycle-Consistent Deformation loss inspired by [48], where the cycle is formed with: a forward deformation from the canonical Gaussians to deformed Gaussians, an anchoring process modifying the deformed Gaussians to the anchored Gaussians, a backward deformation from anchored Gaussians to canonical Gaussians. As a result, we obtain uniformly distributed 3D Gaussians in each time step, and they are aligned with the mesh faces. This makes finding the correspondences between meshes much easier.
@@
We demonstrate high-fidelity mesh reconstruction and tracking in our experiments. A few examples can be found in Fig. 1, where the correspondence is shown as the red curves connecting the vertices across time. Even for challenging geometries (e.g., thin structures like bird wings), our method can still perform reconstruction and achieve much better results than previous approaches, which usually fail to reconstruct the geometry or construct an oversized mesh. We also demonstrate downstream applications, such as texture editing with our extracted mesh and correspondence (Fig. 8), which is much harder to achieve with non-mesh representations. To the best of our knowledge, this is the first framework that recovers high-fidelity meshes with cross-frame correspondences from monocular videos.
@@ 2 Related Work
View Synthesis in a Dynamic Scene. The rise of Neural Radiance Fields (NeRFs) [21] has largely transformed 3D scene reconstruction and novel view sythesis. Its success has also been extended to dynamic scenes in two directions. One direction is to build dynamic NeRFs [4, 7, 16, 41] which models the motion of the scene by extending the radiance field with an extra time dimension or a latent code. To achieve faster rendering speed and more efficient use of memory, recent works perform volume factorization converting the 4D volume into multiple lower dimension planers or tensors [1, 6, 29, 42]. Another direction is to construct an additional deformation field that maps point coordinates in different time frames into a canonical space, where large motion and geometry changes can be captured and learned [23, 27, 24, 35, 5]. The recently proposed 3D Gaussians Splatting [13] extends the volumetric rendering in NeRF by accommodating point clouds. This not only largely improves the speed of neural rendering but also extracts an explicit point cloud structure from images. When applied to dynamic scenes, the same idea of building a deformation field is applied with 3D Gaussians [49, 39, 20]. But instead of an implicit field, the deformation field here explicitly tracks the canonical point clouds over time. Inspired by this line of research, we push forward the reconstruction of explicit geometry representation by using meshes with the tracks of the mesh vertices over time from a monocular video.
@@
Dynamic Mesh Reconstruction. Mesh plays an important role in modern 3D engines, which are widely used for simulation, modeling, and rendering applications. Recovering mesh from static scenes has been extensively studied over the years. Mesh template-based methods [8, 43, 22, 37, 12, 15] optimize the vertex positions from a template to align with the object surface. They usually assume the template is a sphere or a given prior shape, and learn to deform the template into the desired shape to match with the input images. However, when extended to dynamic scenes, these methods often fail because of the topology change during the deformation. Another line of work extract meshed from the implicit field, which can easily handle topology change during deformation. In particular, the mesh can be extracted by identifying and triangulating the zero level sets with methods like Marching Cubes [3], Marching Tetrahedra [34], and Dual Contouring [10]. Recently, differentiable mesh extraction methods have been proposed [17, 2, 28, 30, 31, 38], directly optimizing mesh through implicit filed. Besides static scenes, several works have been studying extracting the surface geometry of deformable objects from dynamic monocular inputs [47, 46, 40, 44, 45, 36, 9], which are more closely related to real-world scenarios. Due to the limited information provided by monocular inputs, these methods either rely on category template information or require strong regularization or data modalities to output satisfying results.
@@ 3 Method
In this section, we present our framework shown in Fig. 3. Given monocular inputs from a dynamic scene, as well as the time label and camera parameters for each input, we reconstruct the dynamic mesh and its appearance, along with the mesh vertex’s motion trajectory. The surface correspondence across times benefits many downstream applications, such as dynamic texture editing, which propagates the editing under a single frame to the rest of the frames.
Our method maintains a set of canonical 3D Gaussians and deformation networks to deform the Gaussians into different times (Sec. 3.1). We use a differentiable Poison Solver and a differentiable Marching Cubes method to reconstruct the oriented 3D Gaussian points into meshes (Sec. 3.2). Our model allows the mesh-to-mesh correspondence. This is constructed by two types of correspondence, including the mesh-to-point correspondence under each time step and the point-to-canonical-point correspondence across all time frames. We propose the Gaussian-Mesh Anchoring to encourage the 1-1 correspondence between the mesh faces and the Gaussians, producing more uniformly distributed anchored Gaussians. The Cycle-Consistent Deformation maintains the points to canonical points correspondence, which includes a backward deformation network that computes the deformation from the anchored Gaussians to the canonical Gaussians (Sec. 3.3). Finally, we combine all the training objectives (Sec. 3.4).
@@ 3.1 Deformable 3D Gaussian Splatting
3D Gaussian Splatting. Recently, 3D Gaussians Splatting (3DGS) [13] has adopted a novel approach based on point cloud rendering, achieving optimal results in novel viewpoint synthesis and scene modeling. In our work, 3DGS provides a fast modeling to explicit geometry and motion, allowing a high-quality and efficient surface reconstruction from the 3D Gaussian point cloud.
@@
Deformable 3D Gaussians. 3DGS assumes static scenes and its performance drops drastically when reconstructing dynamic scenes. Several works [39, 49, 11, 18, 50] have investigated the way to extend the 3DGS to dynamic reconstruction and an efficient way is to maintain a canonical 3D Gaussains set. We define a transformation function $\mathcal{F}$ which transforms the 3D Gaussian $G(x;\mu,r,s,\alpha)$ from one space to another by:
@@
, where $\gamma$ is the positional encoding function to map $x$ and $t$ to higher dimensional Fourier features [32] to enhance detailed reconstruction:
@@ 3.2 Mesh Reconstruction
Surface Extraction and Rendering Pipeline. To apply the isosurface algorithms (such as Marching Cubes [19] or DMTet [30]) and extract the surface geometry from the point-cloud representation of the 3DGS, an indicator function $\chi(x)$ is required over the 3D grid to describe the geometry. Some work [33] query a density grid using 3D Gaussian’s opacity, which does not satisfy the zero derivative condition in the non-boundary space and is computationally inefficient. Instead, we treat 3DGS as an oriented point-cloud and perform a differentiable Poisson solver (DPSR) [26] to obtain the $\chi(x)$. We then apply differentiable Marching Cubes [38] to generate watertight manifold meshes from the obtained value grid. We additionally optimize vertex color for the mesh and perform differentiable rasterization [14] to render the final image.
@@ 3.3 Dynamic Mesh Correspondence
Gaussian-Mesh Anchoring. To build the mesh-to-point correspondence at each time frame, we introduce the Gaussian-Mesh Anchoring. The original 3D Gaussian densification and deformation mechanism leads to uneven spreading of Gaussians in 3D space (Fig. 2), which is undesirable for finding the 1-1 correspondence between mesh and the Gaussians. On the other hand, DPSR discretizes the function values and differential operators in the Poisson equation by assuming the normal vector field $n$ and indicator function $\chi$ are uniformly sampled along each dimension. However, this uniformity is not guaranteed during the above training procedure. To encourage the uniformity of the 3D Gaussian distribution, we use reconstructed mesh to guide the densification and pruning of the 3D Gaussian in the deformed space. Specifically, given a deformed 3D Gaussian set $\{G(x_{i})\}_{i=0}^{N}$ and the set of the face centroids from its reconstructed mesh $\{f_{j}\}_{j=0}^{M}$ and $N\neq M$. For each 3D Gaussian point $x_{i}$ we find its nearest neighbor in the mesh face set:
@@
During the densification and pruning process, as shown in Fig. 4, for face $f_{j}$ that is the nearest neighbor to multiple $K$ Gaussian points, we prune the original $K$ Gaussians and create a new Gaussian by averaging the properties from the $K$ Gaussians:
@@
where $n$ is the total number of faces that have a one-to-one correspondence to the 3D Gaussian points. The obtained anchored Gaussians $G^{\prime}$ are applied in the Cycle-Consistent deformation loss $\mathcal{L}_{cycle}$ which will be introduced in the next paragraph.
$\mathcal{F}_{f}(G,t)=-\mathcal{F}_{d}(G^{\prime},t)$
We use the $L_{1}$ loss between $\mathcal{F}_{f}$ and $-\mathcal{F}_{d}$ as our cycle consistent loss $\mathcal{L}_{cycle}$. This cycle-consistent deformation allows the anchored Gaussian to be deformed back to the canonical space. Specifically, $G^{\prime}$ will be the anchored Gaussians after the Gaussian-Mesh Anchoring, as introduced previously. This cycle-consistency constraint will encourage adjustments to the canonical Gaussians. This process is visualized on the left side of Fig. 3, where the green network is the forward deformation and the blue network is the backward deformation.
3.4 Combining All Training Objectives
We use rendering loss from both mesh rasterization (“Mesh Image” in Fig. 3) and the 3DGS (“Gaussian Image” in Fig. 3) to optimize the mesh geometry and appearance. Given the ground truth image $I_{gt}$ and the Gaussian splatted image $I_{gs}$. The image loss term of 3DGS can be described as:
Same for the rasterized mesh image loss $\mathcal{L}_{mesh}$. We also compute the $L_{1}$ loss on the rasterized mask $\mathcal{L}_{mask}$.
The Laplacian loss term $\mathcal{L}_{lap}$ in Sec. 3.2 helps preserve the mesh surface tessellation and produces smoother surface. The anchoring loss term $\mathcal{L}_{anchor}$ and the cycle-consistency loss term $\mathcal{L}_{cycle}$ described in Sec. 3.3 help the optimization of cross-frame mesh-to-mesh correspondence. The final loss term can be described as:
$\mathcal{L}=$ $\mathcal{L}_{gs}+\mathcal{L}_{mesh}+\mathcal{L}_{mask}$
$+\mathcal{L}_{lap}+\mathcal{L}_{anchor}+\mathcal{L}_{cycle}$

### v3

# Dynamic Gaussians Mesh: Consistent Mesh Reconstruction From Dynamic Scenes

@@ Abstract
Modern 3D engines and graphics pipelines require mesh as a memory-efficient representation, which allows efficient rendering, geometry processing, texture editing, and many other downstream operations. However, it is still highly difficult to obtain high-quality mesh in terms of detailed structure and time consistency from dynamic observations. To this end, we introduce Dynamic Gaussians Mesh (DG-Mesh), a framework to reconstruct a high-fidelity and time-consistent mesh from dynamic input. Our work leverages the recent advancement in 3D Gaussian Splatting to construct the mesh sequence with temporal consistency from dynamic observations. Building on top of this representation, DG-Mesh recovers high-quality meshes from the Gaussian points and can track the mesh vertices over time, which enables applications such as texture editing on dynamic objects. We introduce the Gaussian-Mesh Anchoring, which encourages evenly distributed Gaussians, resulting better mesh reconstruction through mesh-guided densification and pruning on the deformed Gaussians. By applying cycle-consistent deformation between the canonical and the deformed space, we can project the anchored Gaussian back to the canonical space and optimize Gaussians across all time frames. During the evaluation on different datasets, DG-Mesh provides significantly better mesh reconstruction and rendering than baselines.
Figure 1: We propose DG-Mesh, a framework that reconstructs high-fidelity time-consistent mesh for dynamic scenes with complex non-rigid deformations. Given dynamic input and the camera parameters, our method reconstructs the high-quality surface and its appearance, as well as the mesh vertice motion across time frames. Our method can reconstruct mesh with flexible topology change as shown above. Additional results can be found on: https://www.liuisabella.com/DG-Mesh.
Codes and data are publicly available at https://github.com/Isabella98Liu/DG-Mesh.
@@ 1 Introduction
The advent of neural rendering techniques has sparked a surge in research focused on extracting the geometry and motion of dynamic scenes from videos using neural representations. Most studies have concentrated on learning deformable neural radiance fields using volumetric representations (Mildenhall et al., 2021). However, these volumetric models often fall short in terms of memory efficiency and explicit geometric details. The recent introduction of 3D Gaussian Splatting (Kerbl et al., 2023) shifts towards point cloud representations that not only are more memory-efficient but also can provide explicit geometries and superior rendering quality. Further advancements demonstrate how 3D Gaussian Splatting’s explicit geometry can effectively represent dynamic scenes by tracking each point’s movement across frames (Luiten et al., 2023; Wu et al., 2023a; Yang et al., 2023d).
This paper takes a significant step in the realm of explicit geometry and motion estimation for dynamic scenes. We introduce a method to extract high-fidelity meshes and their motions by tracking vertices over time from videos, as illustrated in Figure 3. Meshes, in contrast to volume, offer a more memory-efficient format. The mapping of correspondences on mesh vertices greatly simplifies texture editing and propagation on target objects. Crucially, the dynamic mesh extracted via our method can be seamlessly integrated into a physical simulator, enabling rapid, physics-based rendering adaptable to diverse materials and lighting conditions.
@@
Figure 2: The 3D Gaussian centers before and after the Gaussian-Mesh Anchoring.
@@
We achieve this by decomposing the mesh-to-mesh correspondence into two separate correspondences that are easier to acquire: the mesh-to-points correspondence that maps the mesh faces to the Gaussian points in each time frame; and the points-to-canonical-points correspondence that helps tracking the movement of mesh across time. Our key observation is that the direct training of 3D Gaussians mentioned above will lead to uneven spreading of Gaussians in 3D space (as shown in Figure 2 left). However, this is detrimental for mesh reconstruction performance due to the violation of the Possion solver assumption that points should be evenly distributed; worse still, non-uniform Gaussians make it trickier to track across a long time period. Thus, we propose the Gaussian-Mesh Anchoring procedure during training, which yields uniformly distributed Gaussians at each frame. Concretely, first, we encourage deformed 3D Gaussians in each time step to align with the corresponding mesh surface that has uniform face distribution due to the underlying iso-surface algorithm; and second, we allow adding Gaussians in the world frame at each time step if certain mesh faces are not covered by existing deformed Gaussians, as well as deduplicting Gaussians for mesh faces that are covered multiple times. We call the new set of Gaussians with 1-1 correspondences to mesh faces as anchored Gaussians. As depicted in Figure 2, anchored Gaussians spread more uniformly in 3D.
@@
Finally, we need to update the canonical space to accommodate the added and removed Gaussians in the world frame so that canonical and anchored Gaussians are always consistent. This can be done by introducing a backward deformation network that computes the deformation from the anchored Gaussians to the canonical Gaussians, by which we can inject new Gaussians in the canonical space or remove unnecessary Gaussians. We define a Cycle-Consistent Deformation loss inspired by (Yang et al., 2023b), where the cycle is formed with: a forward deformation from the canonical Gaussians to deformed Gaussians, an anchoring process modifying the deformed Gaussians to the anchored Gaussians, a backward deformation from anchored Gaussians to canonical Gaussians. As a result, we obtain uniformly distributed 3D Gaussians in each time step, and they are aligned with the mesh faces. This makes finding the correspondences between meshes much easier.
@@
We demonstrate high-fidelity mesh reconstruction and tracking in our experiments. A few examples can be found in Figure 1, where the correspondence is shown as the red curves connecting the vertices across time. Even for challenging geometries (e.g., thin structures like bird wings), our method can still perform reconstruction and achieve much better results than previous approaches, which usually fail to reconstruct the geometry or construct an oversized mesh. We also demonstrate downstream applications, such as texture editing with our extracted mesh and correspondence (Figure 26), which is much harder to achieve with non-mesh representations. To the best of our knowledge, this is the first framework that recovers high-fidelity meshes with cross-frame correspondences from dynamic observations.
@@ 2 Related Work
View Synthesis in a Dynamic Scene. The rise of Neural Radiance Fields (NeRFs) (Mildenhall et al., 2021) has largely transformed 3D scene reconstruction and novel view sythesis. Its success has also been extended to dynamic scenes in two directions. One direction is to build dynamic NeRFs (Du et al., 2021; Gao et al., 2021; Li et al., 2021; Xian et al., 2021) which models the motion of the scene by extending the radiance field with an extra time dimension or a latent code. To achieve faster rendering speed and more efficient use of memory, recent works perform volume factorization converting the 4D volume into multiple lower dimension planers or tensors (Cao & Johnson, 2023; Fridovich-Keil et al., 2023; Shao et al., 2023; Xu et al., 2023). Another direction is to construct an additional deformation field that maps point coordinates in different time frames into a canonical space, where large motion and geometry changes can be captured and learned (Park et al., 2021a; Pumarola et al., 2021; Park et al., 2021b; Tretschk et al., 2021; Fang et al., 2022). The recently proposed 3D Gaussians Splatting (Kerbl et al., 2023) extends the volumetric rendering in NeRF by accommodating point clouds. This not only largely improves the speed of neural rendering but also extracts an explicit point cloud structure from images. When applied to dynamic scenes, the same idea of building a deformation field is applied with 3D Gaussians (Yang et al., 2023d; Wu et al., 2023a; Luiten et al., 2023; Li et al., 2024; Yang et al., 2023c). But instead of an implicit field, the deformation field here explicitly tracks the canonical point clouds over time. Inspired by this line of research, we push forward the reconstruction of explicit geometry representation by using meshes with the tracks of the mesh vertices over time from dynamic observations.
@@
Dynamic Mesh Reconstruction. Mesh plays an important role in modern 3D engines, which are widely used for simulation, modeling, and rendering applications. Recovering mesh from static scenes has been extensively studied over the years. Mesh template-based methods (Hanocka et al., 2020; Xue et al., 2023; Pan et al., 2019; Wang et al., 2018; Kanazawa et al., 2018; Li et al., 2020) optimize the vertex positions from a template to align with the object surface. They usually assume the template is a sphere or a given prior shape, and learn to deform the template into the desired shape to match with the input images. However, when extended to dynamic scenes, these methods often fail because of the topology change during the deformation. Another line of work extract meshed from the implicit field, which can easily handle topology change during deformation. In particular, the mesh can be extracted by identifying and triangulating the zero level sets with methods like Marching Cubes (Cubes, 1987), Marching Tetrahedra (Treece et al., 1999), and Dual Contouring (Ju et al., 2002). Recently, differentiable mesh extraction methods have been proposed (Liao et al., 2018; Chen et al., 2022; Remelli et al., 2020; Shen et al., 2021; 2023; Wei et al., 2023), directly optimizing mesh through implicit filed. Besides static scenes, several works have been studying extracting the surface geometry of deformable objects from dynamic monocular inputs (Yang et al., 2023a; 2022; Wu et al., 2023b; Yang et al., 2021a; b; Tulsiani et al., 2020; Johnson et al., 2023), which are more closely related to real-world scenarios. Due to the limited information provided by monocular inputs, these methods either rely on category template information or require strong regularization or data modalities to output satisfying results.
@@ 3 Method
In this section, we present our framework shown in Figure 3. Given inputs from a dynamic scene, as well as the time label and camera parameters for each input, we reconstruct the dynamic mesh and its appearance, along with the mesh vertex’s motion trajectory. The surface correspondence across times benefits many downstream applications, such as dynamic texture editing, which propagates the editing under a single frame to the rest of the frames.
Our method maintains a set of canonical 3D Gaussians and deformation networks to deform the Gaussians into different times (Section 3.1). We use a differentiable Poison Solver and a differentiable Marching Cubes method to reconstruct the oriented 3D Gaussian points into meshes (Section 3.2). Our model allows the mesh-to-mesh correspondence. This is constructed by two types of correspondence, including the mesh-to-point correspondence under each time step and the point-to-canonical-point correspondence across all time frames. We propose the Gaussian-Mesh Anchoring to encourage the 1-1 correspondence between the mesh faces and the Gaussians, producing more uniformly distributed anchored Gaussians. The Cycle-Consistent Deformation maintains the points to canonical points correspondence, which includes a backward deformation network that computes the deformation from the anchored Gaussians to the canonical Gaussians (Section 3.3). Finally, we combine all the training objectives (Section 3.4).
@@ 3.1 Deformable 3D Gaussian Splatting
3D Gaussian Splatting. Recently, 3D Gaussians Splatting (3DGS) (Kerbl et al., 2023) has adopted a novel approach based on point cloud rendering, achieving optimal results in novel viewpoint synthesis and scene modeling. In our work, 3DGS provides a fast modeling to explicit geometry and motion, allowing a high-quality and efficient surface reconstruction from the 3D Gaussian point cloud.
@@
Deformable 3D Gaussians. 3DGS assumes static scenes and its performance drops drastically when reconstructing dynamic scenes. Several works (Wu et al., 2023a; Yang et al., 2023d; Jung et al., 2023; Lin et al., 2023; Zielonka et al., 2023; Li et al., 2024; Yang et al., 2023c) have investigated the way to extend the 3DGS to dynamic reconstruction and an efficient way is to maintain a canonical 3D Gaussains set. We define a transformation function $\mathcal{F}$ which transforms the 3D Gaussian $G(x;\mu,r,s,\alpha)$ from one space to another by:
@@
, where $\gamma$ is the positional encoding function to map $x$ and $t$ to higher dimensional Fourier features (Tancik et al., 2020) to enhance detailed reconstruction:
@@ 3.2 Mesh Reconstruction
Surface Extraction and Rendering Pipeline. To apply the isosurface algorithms (such as Marching Cubes (Lorensen & Cline, 1998) or DMTet (Shen et al., 2021)) and extract the surface geometry from the point-cloud representation of the 3DGS, an indicator function $\chi(x)$ is required over the 3D grid to describe the geometry. Some work (Tang et al., 2023) query a density grid using 3D Gaussian’s opacity, which does not satisfy the gradient condition in the non-boundary space and is computationally inefficient. Instead, we treat 3DGS as an oriented point-cloud and perform a differentiable Poisson solver (DPSR) (Peng et al., 2021) to obtain the $\chi(x)$. We then apply differentiable Marching Cubes (Wei et al., 2023) to generate watertight manifold meshes from the obtained value grid. We additionally optimize vertex color for the mesh and perform differentiable rasterization (Laine et al., 2020) to render the final image.
@@
Figure 5: We compare the mesh reconstruction and rendering results of our method with other baselines on the D-NeRF dataset. For each object, we visualize the results under two different time steps and different viewing angles. Our method outperforms others by producing more smooth and detailed structures.
@@ 3.3 Dynamic Mesh Correspondence
Gaussian-Mesh Anchoring. To build the mesh-to-point correspondence at each time frame, we introduce the Gaussian-Mesh Anchoring. The original 3D Gaussian densification and deformation mechanism leads to uneven spreading of Gaussians in 3D space (Figure 2), which is undesirable for finding the 1-1 correspondence between mesh and the Gaussians. On the other hand, DPSR discretizes the function values and differential operators in the Poisson equation by assuming the normal vector field $n$ and indicator function $\chi$ are uniformly sampled along each dimension. However, this uniformity is not guaranteed during the above training procedure. To encourage the uniformity of the 3D Gaussian distribution, we use reconstructed mesh to guide the densification and pruning of the 3D Gaussian in the deformed space. Specifically, given a deformed 3D Gaussian set $\{G(x_{i})\}_{i=0}^{N}$ and the set of the face centroids from its reconstructed mesh $\{f_{j}\}_{j=0}^{M}$ and $N\neq M$. For each 3D Gaussian point $x_{i}$ we find its nearest neighbor in the mesh face set:
@@
During the densification and pruning process, as shown in Figure 4, for face $f_{j}$ that is the nearest neighbor to multiple $K$ Gaussian points, we prune the original $K$ Gaussians and create a new Gaussian by averaging the properties from the $K$ Gaussians:
@@
where $n$ is the total number of faces that have a one-to-one correspondence to the 3D Gaussian points. A detailed description of the algorithm can be found in Appendix A. The obtained anchored Gaussians $G^{\prime}$ are applied in the Cycle-Consistent deformation loss $\mathcal{L}_{cycle}$ which will be introduced in the next paragraph.
@@
$\mathcal{F}_{f}(G,t)=-\mathcal{F}_{b}(G^{\prime},t)$
We use the $L_{1}$ loss between $\mathcal{F}_{f}$ and $-\mathcal{F}_{d}$ as our cycle consistent loss $\mathcal{L}_{cycle}$. This cycle-consistent deformation allows the anchored Gaussian to be deformed back to the canonical space. Specifically, $G^{\prime}$ will be the anchored Gaussians after the Gaussian-Mesh Anchoring, as introduced previously. This cycle-consistency constraint will encourage adjustments to the canonical Gaussians. This process is visualized on the left side of Figure 3, where the green network is the forward deformation and the blue network is the backward deformation.
3.4 Combining All Training Objectives
We use rendering loss from both mesh rasterization (“Mesh Image” in Figure 3) and the 3DGS (“Gaussian Image” in Figure 3) to optimize the mesh geometry and appearance. Given the ground truth image $I_{gt}$ and the Gaussian splatted image $I_{gs}$. The image loss term of 3DGS can be described as:
Same for the mesh image loss $\mathcal{L}_{mesh}$. We also compute the $L_{1}$ loss on the rasterized mask $\mathcal{L}_{mask}$.
The Laplacian loss term $\mathcal{L}_{lap}$ in Section 3.2 helps preserve the mesh surface tessellation and produces smoother surface. The anchoring loss term $\mathcal{L}_{anchor}$ and the cycle-consistency loss term $\mathcal{L}_{cycle}$ described in Section 3.3 help the optimization of cross-frame mesh-to-mesh correspondence. The final loss term can be described as:
$\mathcal{L}=$ $\mathcal{L}_{gs}+\mathcal{L}_{mesh}+\mathcal{L}_{mask}$
$+\mathcal{L}_{lap}+\mathcal{L}_{anchor}+\mathcal{L}_{cycle}$
Figure 6: We present the mesh reconstruction results of our method and other baselines on the DG-Mesh dataset. Our method delivers better mesh quality and geometry. For objects like the flying bird, other methods struggle to recover the thin wings, while ours successfully reconstructs them.